# atomic orbital API Application Programmer Interface

### 06 Mar 2024 v3

@@ Acknowledgments
We would like to thank Prof. Dr. Trond Saue from the CNRS/Université de Toulouse, France. We acknowledge support from the Research Council of Norway through its Centres of Excellence scheme (262695), through the FRIPRO grant ReMRChem (324590), and from NOTUR – The Norwegian Metacenter for Computational Science through grant of computer time (nn4654k).
@@
We use the CRediT taxonomy of contributor roles.[57, 58] The “Investigation” role also includes the “Methodology”, “Software”, and “Validation” roles. The “Analysis” role also includes the “Formal analysis” and “Visualization” roles. The “Funding acquisition” role also includes the “Resources” role. We visualize contributor roles in the following authorship attribution matrix, as suggested in Ref. 59.

### 27 Feb 2026 v4

@@ Acknowledgments
We would like to thank Prof. Dr. Trond Saue from the CNRS/Université de Toulouse, France for fruitful discussions on the topic. We acknowledge support from the Research Council of Norway through its Centres of Excellence scheme (262695), through the FRIPRO grant ReMRChem (324590), and from NOTUR – The Norwegian Metacenter for Computational Science through grant of computer time (nn14654k).
@@
We use the CRediT taxonomy of contributor roles.[Allen2014-rd, Brand2015-qc] The “Investigation” role also includes the “Methodology”, “Software”, and “Validation” roles. The “Analysis” role also includes the “Formal analysis” and “Visualization” roles. The “Funding acquisition” role also includes the “Resources” role. We visualize contributor roles in the following authorship attribution matrix, as suggested in Ref. authorship-attribution.